OpenStudy (anonymous):

Find general solution to differential equation. 16y'' +8y' +y = 0

OpenStudy (anonymous):

Characteristic equation is 16r^2 + 8r +1 = 0

OpenStudy (anonymous):

I thought to use quadratic to solve...

OpenStudy (anonymous):

sorry, actually it is factorable.

OpenStudy (rational):

yep

OpenStudy (anonymous):

SO is general solution then y = C1e^-t/4 + C2e^-t/4 ?

OpenStudy (rational):

nope, you need two independent solutions

OpenStudy (rational):

you need to find another solution which doesnt look like e^-t/4

OpenStudy (anonymous):

ok.....any tips where to start.

OpenStudy (rational):

say the second independent solution is \(u(t) e^{\frac{-t}{4}}\)

OpenStudy (rational):

\(y = u* e^{\frac{-t}{4}}\) \(y' = ?\) \(y'' = ?\)

OpenStudy (anonymous):

just a sec..

OpenStudy (anonymous):

y' = v'e^-t/4 - 1/4 *e^-t/4 * v y'' = v''e^-t/4 - 1/4 *e^/t/4 *v' + 1/8 *e^-1/4 * v -1/4 *e^-t/4 v'

OpenStudy (rational):

yes combine v, v' and v'' terms and write it in a readable form lol

OpenStudy (anonymous):

you mean combine the two v' terms I am assuming.

OpenStudy (rational):

Yep \(y = u* e^{\frac{-t}{4}} \) \(y' = u' e^{\frac{-t}{4} } - u\frac{1}{4} e^{\frac{-t}{4}} \) \(y'' = u'' e^{\frac{-t}{4} } - u'\frac{1}{4}e^{\frac{-1}{4}} - u' \frac{1}{4} e^{\frac{-t}{4}} + u \frac{1}{16} e^{\frac{-t}{4}} \) \(~~~~ = u'' e^{\frac{-t}{4} } - u'\frac{1}{2}e^{\frac{-1}{4}} + u \frac{1}{16} e^{\frac{-t}{4}} \)

OpenStudy (rational):

see if that looks okay

OpenStudy (anonymous):

same thing I wrote...thanks I got it from here.

OpenStudy (rational):

you should get \(u(t) = t\)

OpenStudy (anonymous):

let me post a photo of what I have....

OpenStudy (rational):

okie

OpenStudy (anonymous):

Not sure what to do at the end ); So is it because it is one repeated soltuion to original characteristic that we have to find a unique seperate solution this way?

OpenStudy (anonymous):

If I remember right, if you have something like (m+1)^2 the general solution is \[\huge y = C_1 e^{-m}+ C_2Xe^{-m}\] or something like that. What I mean is if you have a square, you just put something like an X with a higher degree. I may be remembering wrong though

OpenStudy (anonymous):

Sorry photo almost too small to read.

OpenStudy (anonymous):

Thats what I thought for a minute also but not sure....

OpenStudy (rational):

adding up should give u : \(16 e^{\frac{-t}{4}} u'' = 0\)

OpenStudy (rational):

solving \(u'' = 0\) gives \(u = t\) \(\implies \) second solution is \(u*e^{\frac{-t}{4}} = te^{\frac{-t}{4}}\)

OpenStudy (anonymous):

@jasonjohnson86 i think that was the same potato as what rational has just said in the end

OpenStudy (anonymous):

right....I dont think I need to run through all that for this solution though.....

OpenStudy (anonymous):

It's just a repeated factor...not reduction of order right?

OpenStudy (rational):

you dont need to, but u need to knw why you're multiplying \(t\)

OpenStudy (anonymous):

for non-BS Math students, I personally think it's sufficient to just know that you multiply a variable of a higher degree for repeated factors =))

OpenStudy (rational):

yes, if today is your last day in differential equations course